OpenStudy (anonymous):

This a question from a test. Its not the exact question but just the short notes I took of it. All help is appreciated. this is a MATRIX ques. im given: AX=C A=[a c] 4x+2y=7 [b d] 5x-6y=9 the question tells me to find the value of a-b+c+d

OpenStudy (zzr0ck3r):

this is very hard to see what you are asking the equations 4x+2y=7, and 5x-6y=9 can be modlled by the matrix equation Ax=c where A = [4,2;5,-6], x = [x:y] and c=[7;9] giving a-b+c+d = 4-2+5-(-6) = 13

OpenStudy (zzr0ck3r):

modeled*

OpenStudy (zzr0ck3r):

err 4-2+5-6 = 1

OpenStudy (zzr0ck3r):

I hope this is what you are asking

OpenStudy (cp9454):

\[A= \left[\begin{matrix}4 & 2 \\5 & -6\end{matrix}\right], X= \left(\begin{matrix}x \\ y\end{matrix}\right), and C = \left(\begin{matrix}7 \\ 9\end{matrix}\right)\]

OpenStudy (cp9454):

solve AX=B

OpenStudy (zzr0ck3r):

they dont want X they want a-b+c+d

OpenStudy (anonymous):

thank you both for your help, but im still confused.

OpenStudy (zzr0ck3r):

the matrix equation Ax = c, where the coefficients are as you listed a,b,c,d is just another way of representing the system of equations ax+by=c_1, and cx+dy=c_2 here a = 4, b = 2 , c = 5, and d = -6, c_1 = 7 and c_2 = 9

OpenStudy (anonymous):

okay, that makes more sence. thank you.

OpenStudy (zzr0ck3r):

if I give you 2x+3y+5z=5 3x+7y-2z = 4 7x+2y+4y = 1 you would have the equation \(Ax=c\), where \(A= \left[\begin{matrix}2 & 3 & 5\\3 & 7 & -2 \\7 & 2 & 4\end{matrix}\right], x= \left(\begin{matrix}x \\ y\end{matrix}\right), \text{ and } {c} = \left(\begin{matrix}5 \\ 4 \\ 1\end{matrix}\right)\)